# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ 4.3.1 Peaked spectrum
We should also discuss an observational constraint on $f^{ss}_{{\rm NL},0}(k)$. So far, the CMB temperature bispectrum $B_{TTT}$ (i.e., a non-linearity parameter for a scalar non-Gaussianity) has been roughly constrained as $B_{TTT}/P_{\zeta}^{2}\lesssim 10$ [108]. Since both scalar and tensor perturbations generate the CMB temperature fluctuations (i.e., $\mathcal{B}_{shh}$ contributes to $B_{TTT}$), the above constraint naively applies to the STT bispectrum as
$\mathcal{B}_{shh}|_{k_{\rm CMB}}\sim\mathcal{O}\biggl{(}\frac{b_{1}}{\mathcal{G}_{T}},\frac{b_{2}}{\mathcal{F}_{T}}\biggr{)}\times P_{\zeta}|_{k_{\rm CMB}}P_{h}|_{k_{\rm CMB}}\lesssim 10P^{2}_{\zeta}|_{k_{\rm CMB}},$ (4.36)
which yields the following constraint
$\mathcal{O}\biggl{(}\frac{b_{1}}{\mathcal{G}_{T}},\frac{b_{2}}{\mathcal{F}_{T}}\biggr{)}\lesssim 10^{3}\,\left(\frac{r|_{k_{\rm CMB}}}{10^{-2}}\right)^{-1}.$ (4.37)
@@
For $r|_{k_{\rm CMB}}=10^{-2}$, we obtain a stronger constraint on $f^{ss}_{{\rm NL},0}(k)$ (i.e., $f^{ss}_{{\rm NL},0}(k)\lesssim 10^{3}$) than the theoretical one in Eq. (4.35). Finally, we can estimate the possible maximum values of $C^{\rm GW}_{\ell}$ and $C^{\rm GW-T}_{\ell}$. Since we have assumed the Bunch-Davies scalar mode, we have $\mathcal{P}_{\zeta}\simeq 2\times 10^{-9}$. Also, the upper bound on $f^{ss}_{{\rm NL},0}(k)$ is evaluated as $10^{3}$ as discussed above. As a result, we can find that $C^{\rm GW}_{\ell}$ and $C^{\rm GW-T}_{\ell}$ can reach $10^{-4}$ and $10^{-7}$, respectively. The allowed values of the correlation functions are shown in Fig. 2.
@@
Figure 2: Left: The allowed values of $C_{\ell}^{\rm GW}$. Right: The allowed values of $C_{\ell}^{\rm GW-T}$.
Let us mention the possibility of obtaining a large value of $b_{1,2}$ in the framework of slow-roll inflation. So far, we have shown that $f^{ss}_{{\rm NL},0}(k)$ can be enhanced by a factor $-k\eta_{0}\,(\lesssim 10)$ due to the consequence of the non-Bunch-Davies effects. The largest enhancement of $f^{ss}_{{\rm NL},0}(k)$ can thus arise from large values of the coupling functions of the cubic interactions $b_{1,2}$. Thus, one might expect that similar enhancements could be predicted from inflation with the Bunch-Davies state. However, in that case, the squeezed bispectrum is expected to satisfy the so-called consistency relation between the spectral index $n_{t}$ and the amplitude of the squeezed bispectrum [37, 59]. To show it explicitly in the Horndeski theory, one would need to take into account slow-roll corrections to the scale factor and the mode functions as has been studied in the context of the scalar non-Gaussianity in Ref. [109]. Without taking those into account, one can show at least that the squeezed bispectrum with the Bunch-Davies state is suppressed by the slow-roll parameters. Now the terms in the case of the Bunch-Davies state correspond to the ones irrelevant with $k\eta_{0}$ in Eq. (4.22). The explicit forms of $b_{1}$ and $b_{2}$ are obtained as
$b_{1}$ $=\frac{3\mathcal{G}_{T}}{8}\biggl{\{}1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}+\frac{\mathcal{G}_{T}}{3}\biggl{[}-\frac{\dot{\mathcal{F}}_{T}}{\mathcal{F}_{T}^{2}}\frac{\mathcal{G}_{T}}{\Theta}+\frac{1}{\mathcal{F}_{T}}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mathcal{G}_{T}}{\Theta}\biggr{)}\biggr{]}\biggr{\}},$ (4.38)
$b_{2}$ $=\frac{1}{8}\frac{H\mathcal{G}_{T}^{2}}{\Theta}+\frac{1}{8}\frac{\mathcal{G}_{T}}{\Theta}\dot{\mathcal{G}}_{T}+\frac{\mathcal{G}_{T}}{8}\frac{{\rm d}}{{\rm d}t}\biggl{(}\frac{\mathcal{G}_{T}}{\Theta}\biggr{)}-\frac{\mathcal{F}_{T}}{8}.$ (4.39)
@@
$\frac{b_{1}}{\mathcal{G}_{T}}+\frac{3b_{2}}{\mathcal{F}_{T}}$ $=\frac{\sigma}{8c_{h}^{2}}\biggl{(}2\epsilon-f_{T}+3g_{T}+\frac{2\dot{\sigma}}{H}\biggr{)},$ (4.40)
with $\sigma:=H\mathcal{G}_{T}/\Theta$. Note that $\sigma=1$ in the k-essence theory. The above is nothing but the combination appearing in the case of the Bunch-Davies state and is suppressed by the slow-roll parameters. However, for different combinations (e.g., $b_{1}/\mathcal{G}_{T}+b_{2}/\mathcal{F}_{T}$), there are terms that are not proportional to the slow-roll parameters. To see more details about the enhancement of $f^{ss}_{{\rm NL},0}(k)$, let us express our $f^{ss}_{{\rm NL},0}(k)$ in terms of $f_{\rm NL,BD}$ which is the result in the case of the Bunch-Davies state as
$f^{ss}_{{\rm NL},0}(k)=f_{\rm NL,BD}-2c_{h}k\eta_{0}\mathcal{A}(k,k_{*})\biggl{(}f_{\rm NL,BD}-\sqrt{\pi}\frac{b_{2}}{\mathcal{F}_{T}}\frac{1+2c_{h}^{2}k^{2}\eta_{0}^{2}}{c_{h}^{2}k^{2}\eta_{0}^{2}}\biggr{)},$ (4.41)
@@
$f_{\rm NL,BD}:=8\sqrt{\pi}\biggl{(}\frac{b_{1}}{\mathcal{G}_{T}}+\frac{3b_{2}}{\mathcal{F}_{T}}\biggr{)}.$ (4.42)
@@
$r|_{k_{\rm CMB}}\simeq 16\frac{\mathcal{F}_{S}}{\mathcal{F}_{T}}\frac{c_{s}}{c_{h}}.$ (4.43)
@@
$f^{ss}_{{\rm NL},0}(k)=f_{\rm NL,BD}-2c_{h}k\eta_{0}\mathcal{A}(k,k_{*})\biggl{(}f_{\rm NL,BD}-\frac{\sqrt{\pi}}{128}\frac{c_{h}}{c_{s}}r|_{k_{\rm CMB}}\frac{1+2c_{h}^{2}k^{2}\eta_{0}^{2}}{c_{h}^{2}k^{2}\eta_{0}^{2}}\biggr{)}.$ (4.44)
@@
$|f^{ss}_{{\rm NL},0}(k)-f_{\rm NL,BD}|\leq\mathcal{O}\biggl{(}\frac{c_{h}}{c_{s}}r|_{k_{\rm CMB}}\biggr{)}\leq\mathcal{O}\biggl{(}10^{-2}\frac{c_{h}}{c_{s}}\biggr{)},$ (4.45)
where we assumed $r|_{k_{\rm CMB}}\sim 10^{-2}$ in the last step, which indicates that one needs a small $c_{s}$. This is dangerous at least within the k-essence theory where a small $c_{s}$ always indicates a large scalar non-Gaussianity [38]. However, this is not the case in more general models of inflation, and in the Horndeski theory, a small $c_{s}$ does not necessarily mean a large scalar non-Gaussianity. Therefore, we conclude that inflation with the non-Bunch-Davies states can potentially enhance $f^{ss}_{{\rm NL},0}(k)$ up to $10^{3}$ and the SGWB anisotropies $C_{\ell}^{\rm GW}$ and $C_{\ell}^{\rm GW-T}$ up to $10^{-4}$ and $10^{-7}$, respectively.
Before moving to the scale-invariant case, let us compare the amplitudes of the SGWB anisotropies from inflation with the non-Bunch-Davies state to those originating from the propagation of the GWs at a late time. On large scales, the dominant contributions to the anisotropies originate from the SW effect as in the CMB fluctuations, the auto- and cross-correlation functions originating from which are, respectively, of the form, [29]
$C_{\ell}^{\rm GW,SW}$ $=\frac{128\pi}{9\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 3\times 10^{-9}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right),$ (4.46)
$C_{\ell}^{\rm GW-T,SW}$ $=\frac{16\pi}{15\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 2\times 10^{-10}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right).$ (4.47)
@@
4.3.2 Nearly-scale-invariant spectrum
In the case of the $k$-independent Bogoliubov coefficients, [^4] In the scale-invariant case, the contribution which is proportional to $\hat{n}\cdot\hat{q}$ in the scalar-tensor-tensor bispectrum is zero in the squeezed limit. we have
${\rm Re}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]=1,$ (4.48)
@@ 4.3.2 Nearly-scale-invariant spectrum
$\qquad\qquad\qquad=-2B.$ (4.49)
@@
$f^{ss}_{{\rm NL},0}(k)=2\sqrt{\pi}\biggl{\{}\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)+2B\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\biggr{\}}.$ (4.50)
Since the above $f^{ss}_{{\rm NL},0}(k)$ is independent of $q$, calculations of the correlation functions can be similarly performed as in the previous case with peaked power spectrum, and $C_{\ell}^{\rm GW}$ and $C_{\ell}^{\rm GW-T}$ are given by Eqs. (4.23) and (4.29), respectively.
As for the constraints on the non-linearity parameters, the constraints on the coupling functions of the cubic interactions $b_{1,2}$ are also expected to be roughly given by Eqs. (4.30), (4.31) and (4.37). On the other hand, the theoretical backreaction constraint reads
$c_{h}B^{2}M_{\rm Pl}^{4}\lesssim M_{\rm Pl}^{2}H_{\rm inf}^{2},$ (4.51)
where we have introduced a cut-off scale $\Lambda=M_{\rm Pl}$ for the physical momentum ($k/a(\eta)$) at $\eta=\eta_{0}$. We thus have $B\lesssim 10^{-5}$ for $H_{\rm inf}=10^{13}~{}{\rm GeV}$ and $c_{h}\simeq 1$, and
$-k\eta_{0}=\frac{k}{a(\eta_{0})H_{\rm inf}}=\frac{k_{\rm max}}{a(\eta_{0})H_{\rm inf}}\frac{k}{k_{\rm max}}=\frac{M_{\rm Pl}}{H_{\rm inf}}\frac{k}{k_{\rm max}}\sim 10^{5}\frac{k}{k_{\rm max}}\leq 10^{5},$ (4.52)
which indicates
$(-k\eta_{0})B\lesssim 1.$ (4.53)
Here, one has $(-k\eta_{0})B\sim 1$ only for the case of $k_{\rm max}=k$, i.e., $k/a(\eta_{0})\sim M_{\rm Pl}$. Thus, in the scale-invariant case the amplitude of SGWB anisotropies is found to be determined by the cubic coupling functions, $b_{1,2}$ even with taking the non-Bunch-Davies vacuum states into account. Nevertheless, contrary to the Bunch-Davies vacuum, we can expect the violation of Maldacena’s consistency relation as discussed above, and hence the maximum amplitudes of the anisotropies in the scale-invariant spectrum case are the same as those in the peaked spectrum case.

### v3

@@ 4.3.1 Peaked spectrum
Here, we have chosen the parameters in the Bogoliubov coefficients $A$ and $\Delta$ in such a way that the SGWBs can be detected with PTA experiments. As a result, the backreaction constraint has put a constraint on only $k_{*}\eta_{0}$ as in Eq. (116) that does not depend on the choice of the cutoff scale. We should also discuss an observational constraint on $f^{ss}_{{\rm NL},0}(k)$. So far, the CMB temperature bispectrum $B_{TTT}$ (i.e. a non-linearity parameter for a scalar non-Gaussianity) has been roughly constrained as $B_{TTT}/P_{\zeta}^{2}\lesssim 10$ Planck:2019kim . Since both scalar and tensor perturbations generate the CMB temperature fluctuations (i.e. $\mathcal{B}_{shh}$ contributes to $B_{TTT}$), the above constraint naively applies to the STT bispectrum as
$\mathcal{B}_{shh}|_{k_{\rm CMB}}\sim\mathcal{O}\biggl(\frac{b_{1}}{\mathcal{G}_{T}},\frac{b_{2}}{\mathcal{F}_{T}}\biggr)\times P_{\zeta}|_{k_{\rm CMB}}P_{h}|_{k_{\rm CMB}}\lesssim 10P^{2}_{\zeta}|_{k_{\rm CMB}},$ (118)
which yields the following constraint:
$\mathcal{O}\biggl(\frac{b_{1}}{\mathcal{G}_{T}},\frac{b_{2}}{\mathcal{F}_{T}}\biggr)\lesssim 10^{3}\,\left(\frac{r|_{k_{\rm CMB}}}{10^{-2}}\right)^{-1}.$ (119)
@@
For $r|_{k_{\rm CMB}}=10^{-2}$, we obtain a stronger constraint on $f^{ss}_{{\rm NL},0}(k)$ (i.e. $f^{ss}_{{\rm NL},0}(k)\lesssim 10^{4}$) than the theoretical one in Eq. (117). Finally, we can estimate the maximum possible values of $C^{\rm GW}_{\ell}$ and $C^{\rm GW-T}_{\ell}$. Since we have assumed the Bunch-Davies scalar mode, we have $\mathcal{P}_{\zeta}\simeq 2\times 10^{-9}$. Also, the upper bound on $f^{ss}_{{\rm NL},0}(k)$ is evaluated as $10^{4}$ as discussed above. As a result, we can find that $C^{\rm GW}_{\ell}$ and $C^{\rm GW-T}_{\ell}$ can reach $10^{-2}$ and $10^{-6}$, respectively. The allowed values of the correlation functions are shown in Fig. 2.
@@
Figure 2: Left: The allowed values of $C_{\ell}^{\rm GW}$ for the peaked spectrum. Right: The allowed values of $C_{\ell}^{\rm GW-T}$ for the peaked spectrum.
Let us mention the possibility of obtaining a large value of $b_{1,2}$ in the framework of slow-roll inflation. So far, we have shown that $f^{ss}_{{\rm NL},0}(k)$ can be enhanced by a factor $k\eta_{0}\,(\lesssim 10)$ due to the consequence of the non-Bunch-Davies effects. The largest enhancement of $f^{ss}_{{\rm NL},0}(k)$ can thus arise from large values of the coupling functions of the cubic interactions $b_{1,2}$. Thus, one might expect that similar enhancements could be predicted from inflation with the Bunch-Davies state. However, in that case, the squeezed bispectrum is expected to satisfy the so-called consistency relation between the spectral index $n_{t}$ and the amplitude of the squeezed bispectrum Maldacena:2002vr ; Kundu:2013gha . To show it explicitly in the Horndeski theory, one would need to take into account slow-roll corrections to the scale factor and the mode functions, as has been studied in the context of the scalar non-Gaussianity in Ref. DeFelice:2013ar . Without taking those into account, one can show at least that the squeezed bispectrum with the Bunch-Davies state is suppressed by the slow-roll parameters. Now the terms in the case of the Bunch-Davies state correspond to the ones irrelevant to $k\eta_{0}$ in Eq. (104). The explicit forms of $b_{1}$ and $b_{2}$ are obtained as
$b_{1}$ $=\frac{3\mathcal{G}_{T}}{8}\biggl\{1-\frac{H\mathcal{G}_{T}^{2}}{\Theta\mathcal{F}_{T}}+\frac{\mathcal{G}_{T}}{3}\biggl[-\frac{\dot{\mathcal{F}}_{T}}{\mathcal{F}_{T}^{2}}\frac{\mathcal{G}_{T}}{\Theta}+\frac{1}{\mathcal{F}_{T}}\frac{{\rm d}}{{\rm d}t}\biggl(\frac{\mathcal{G}_{T}}{\Theta}\biggr)\biggr]\biggr\},$ (120)
$b_{2}$ $=\frac{1}{8}\frac{H\mathcal{G}_{T}^{2}}{\Theta}+\frac{1}{8}\frac{\mathcal{G}_{T}}{\Theta}\dot{\mathcal{G}}_{T}+\frac{\mathcal{G}_{T}}{8}\frac{{\rm d}}{{\rm d}t}\biggl(\frac{\mathcal{G}_{T}}{\Theta}\biggr)-\frac{\mathcal{F}_{T}}{8}.$ (121)
@@
$\frac{b_{1}}{\mathcal{G}_{T}}+\frac{3b_{2}}{\mathcal{F}_{T}}$ $=\frac{\sigma}{8c_{h}^{2}}\biggl(2\epsilon-f_{T}+3g_{T}+\frac{2\dot{\sigma}}{H}\biggr),$ (122)
with $\sigma:=H\mathcal{G}_{T}/\Theta$. Note that $\sigma=1$ in the k-essence theory. The above is nothing but the combination appearing in the case of the Bunch-Davies state and is suppressed by the slow-roll parameters. However, for different combinations (e.g. $b_{1}/\mathcal{G}_{T}+b_{2}/\mathcal{F}_{T}$), there are terms that are not proportional to the slow-roll parameters. To see more details about the enhancement of $f^{ss}_{{\rm NL},0}(k)$, let us express our $f^{ss}_{{\rm NL},0}(k)$ in terms of $f_{\rm NL,BD}$ which is the result in the case of the Bunch-Davies state as
$f^{ss}_{{\rm NL},0}(k)=f_{\rm NL,BD}-2c_{h}k\eta_{0}\mathcal{A}(k,k_{*})\biggl(f_{\rm NL,BD}-\sqrt{\pi}\frac{b_{2}}{\mathcal{F}_{T}}\frac{1+2c_{h}^{2}k^{2}\eta_{0}^{2}}{c_{h}^{2}k^{2}\eta_{0}^{2}}\biggr),$ (123)
@@
$f_{\rm NL,BD}:=8\sqrt{\pi}\biggl(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{3b_{2}}{\mathcal{F}_{T}}\biggr).$ (124)
@@
$r|_{k_{\rm CMB}}\simeq 16\frac{\mathcal{F}_{S}}{\mathcal{F}_{T}}\frac{c_{s}}{c_{h}}.$ (125)
@@
$f^{ss}_{{\rm NL},0}(k)=f_{\rm NL,BD}-2c_{h}k\eta_{0}\mathcal{A}(k,k_{*})\biggl(f_{\rm NL,BD}-\frac{\sqrt{\pi}}{128}\frac{c_{h}}{c_{s}}r|_{k_{\rm CMB}}\frac{1+2c_{h}^{2}k^{2}\eta_{0}^{2}}{c_{h}^{2}k^{2}\eta_{0}^{2}}\biggr).$ (126)
@@
$|f^{ss}_{{\rm NL},0}(k)-f_{\rm NL,BD}|\leq\mathcal{O}\biggl(\frac{c_{h}}{c_{s}}r|_{k_{\rm CMB}}\biggr)\leq\mathcal{O}\biggl(10^{-2}\frac{c_{h}}{c_{s}}\biggr),$ (127)
where we assumed $r|_{k_{\rm CMB}}\sim 10^{-2}$ in the last step, which indicates that one needs a small $c_{s}$. This is dangerous at least within the k-essence theory where a small $c_{s}$ always indicates a large scalar non-Gaussianity Chen:2006nt . As will be given in Appendix D, both the maximum enhancement of the SGWB anisotropies and the generation of a small scalar non-Gaussianity are simultaneously achievable in the full Horndeski theory i.e. in the presence of the $G_{4}/G_{5}$ terms. Therefore, we conclude that inflation with the non-Bunch-Davies states in the full Horndeski action can potentially enhance $f^{ss}_{{\rm NL},0}(k)$ up to $10^{4}$ and the SGWB anisotropies $C_{\ell}^{\rm GW}$ and $C_{\ell}^{\rm GW-T}$ up to $10^{-2}$ and $10^{-6}$, respectively.
Before moving to the scale-invariant case, let us compare the amplitudes of the SGWB anisotropies from inflation with the non-Bunch-Davies states to those originating from the propagation of the GWs at a late time. On large scales, the dominant contributions to the anisotropies originate from the SW effect as in the CMB fluctuations; the auto- and cross-correlation functions originating from which are, respectively, of the form Dimastrogiovanni:2021mfs ,
$C_{\ell}^{\rm GW,SW}$ $=\frac{128\pi}{9\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 3\times 10^{-9}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right),$ (128)
$C_{\ell}^{\rm GW-T,SW}$ $=\frac{16\pi}{15\ell(\ell+1)}\mathcal{P}_{\zeta}\simeq 2\times 10^{-10}\left[\frac{30}{\ell(\ell+1)}\right]\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right).$ (129)
@@
4.3.2 Nearly scale-invariant spectrum
In the case of the $k$-independent Bogoliubov coefficients, [^4] In the scale-invariant case, the contribution which is proportional to $\hat{n}\cdot\hat{q}$ in the STT bispectrum is zero in the squeezed limit. we have
${\rm Re}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]=1,$ (130)
@@ 4.3.2 Nearly scale-invariant spectrum
$\qquad\qquad\qquad=-2B.$ (131)
@@
$f^{ss}_{{\rm NL},0}(k)=2\sqrt{\pi}\biggl\{\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)+2B\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\biggr\}.$ (132)
Since the above $f^{ss}_{{\rm NL},0}(k)$ is independent of $q$, calculations of the correlation functions can be similarly performed as in the previous case with peaked power spectrum, and $C_{\ell}^{\rm GW}$ and $C_{\ell}^{\rm GW-T}$ are given by Eqs. (105) and (111), respectively.
As for the constraints on the non-linearity parameters, the constraints on the coupling functions of the cubic interactions $b_{1,2}$ are also expected to be roughly given by Eqs. (112), (113), and (119). On the other hand, the theoretical backreaction constraint reads
$c_{h}B^{2}\Lambda^{4}\lesssim M_{\rm Pl}^{2}H_{\rm inf}^{2},$ (133)
where we have imposed $k/a(\eta_{0})\leq\Lambda$. Note here that the backreaction constraint becomes weaker for a lower cutoff scale, which is in contrast to the peaked-spectrum case in which the backreaction constraint is insensitive to the choice of the cutoff scale. In particular, the parameter $B$ can increase up to $\mathcal{O}(10^{5})$ for the cutoff scale given by Eq. (35). By combining the above with Eq. (43), Eq. (132), and $c_{h}=\mathcal{O}(1)$, we obtain
$f^{ss}_{\rm NL,0}(k)\sim B(-k\eta_{0})\mathcal{O}\left(\frac{b_{1}}{\mathcal{G}_{T}},\frac{b_{2}}{\mathcal{F}_{T}}\right)\leq\frac{M_{\rm Pl}}{\Lambda}\mathcal{O}(10^{3}),$ (134)
which holds for any $\Lambda$. One can find from the above that as the cutoff scale is lower (i.e. the backreaction constraint is weaker), the maximum possible magnitude of $f^{ss}_{\rm NL,0}$ becomes greater. In particular, for the cutoff scale given by Eq. (35), we have $f^{ss}_{\rm NL,0}(k)\leq\mathcal{O}(10^{7})$ where $f^{ss}_{\rm NL,0}(k)$ can be $\mathcal{O}(10^{7})$ for the case of $k/a(\eta_{0})\sim\Lambda=(\sqrt{\epsilon}M_{\rm Pl}H_{\rm inf}^{2})^{1/3}$. As a result, $C^{\rm GW}_{\ell}$ and $C^{\rm GW-T}_{\ell}$ can reach $10^{4}$ and $10^{-3}$, respectively. In fact, in this case, since the parameter $B$ can take $O(10^{5})$, it can be considered that the SGWB anisotropies are enhanced by the non-Bunch-Davies states. Note that the violation of the non-Gaussianity consistency relation still plays an important role in enhancing the anisotropies up to the maximum possible amplitudes.
The allowed values of the correlation functions are shown in Fig. 3. The amplitudes of GW anisotropies can be greater in the scale-invariant case than in the peaked-spectrum case in which the maximum value of $f^{ss}_{\rm NL,0}(k)$ is $\mathcal{O}(10^{4})$.
Refer to caption Refer to caption
Figure 3: Left: The allowed values of $C_{\ell}^{\rm GW}$ for the scale-invariant spectrum. Right: The allowed values of $C_{\ell}^{\rm GW-T}$ for the scale-invariant spectrum.